OpenStudy (anonymous):

Use a transformation to solve the equation. d – (–17) = –14 A. d = –31 B. d = –3 C. d = 3 D. d = 31

OpenStudy (anonymous):

ill give metal and become fan

OpenStudy (anonymous):

It is A

OpenStudy (anonymous):

and i think its a

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

Np

OpenStudy (anonymous):

can u help me on more

OpenStudy (anonymous):

Yeah sure

OpenStudy (anonymous):

What is the solution to the equation? c + (–2) = –0.7 c =

OpenStudy (anonymous):

there

OpenStudy (anonymous):

1.3

OpenStudy (anonymous):

Use transformations to solve the equation. y – 0.6 = –1.8 A. y = –2.4 B. y = –1.2 C. y = 1.2 D. y = 2.4

OpenStudy (anonymous):

b

OpenStudy (anonymous):

?

OpenStudy (anonymous):

@Desireelover r u there

OpenStudy (anonymous):

@DenishaH

OpenStudy (anonymous):

Yes I am

OpenStudy (anonymous):

@DenishaH i think its b

OpenStudy (anonymous):

ill brb

OpenStudy (anonymous):

You're correct :D

OpenStudy (anonymous):

im back

OpenStudy (anonymous):

i need help A square is always an example of a

OpenStudy (anonymous):

@JessGirl1093 can u help

OpenStudy (anonymous):

with what

OpenStudy (anonymous):

the qustion

OpenStudy (anonymous):

What is the solution to the equation? y – 0.7 = –2.2 y =

OpenStudy (anonymous):

after this i just need 1 more

OpenStudy (anonymous):

if u help me first my sis wants to know A square is always an example of a

OpenStudy (anonymous):

i think so not sure

OpenStudy (anonymous):

waiting..................................

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

a parallelogram, a rhombus, a rectangle, a regular polygon

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

np now can u help me

OpenStudy (anonymous):

@paki

OpenStudy (anonymous):

@rosy_love

OpenStudy (anonymous):

@Desireelover

OpenStudy (anonymous):

What is the solution to the equation? y – 0.7 = –2.2 y =

OpenStudy (anonymous):

can u help

OpenStudy (anonymous):

2.9

OpenStudy (anonymous):

thx can u help on 1 more

OpenStudy (anonymous):

Use a transformation to solve the equation. t – (–16.5) = –13.4 A. t = –29.9 B. t = –3.1 C. t = 3.1 D. t = 29.9

OpenStudy (anonymous):

i think its d

OpenStudy (anonymous):

It is A

OpenStudy (anonymous):

thx do u wanna no wt i got

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

yessss i got 80'/' thx

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

byeee